# The Klein-Gordon equation on asymptotically Minkowski spacetimes: causal propagators

### v1

@@ 5.2. Further elliptic regularity near $C$
This definition of $F_{\psi}$ makes sense as a bounded operator on $L^{2}$ by continuous functional calculus since $(D_{t}^{2}+H_{V_{0}}+E)^{-1}P_{V_{0}}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ is a bounded symmetric operator. Note that $f_{\psi}$ is not compactly supported, so the results of Section 5 do not apply directly.
If we knew that $E_{\psi}\in{}^{3\mathrm{sc}}\Psi^{-2,0}$ then we would immediately have estimates. That is the purpose of the following lemma.
Lemma 5.12.
Let $\psi\in C_{c}^{\infty}(\mathbb{R})$ with $\psi(s)=1$ for $|s|<\delta$. Then there is $\chi_{\mathrm{NP}}\in C^{\infty}(X)$ supported sufficiently close to $\mathrm{NP}$ and
$E_{\psi}\in{}^{3\mathrm{sc}}\Psi^{-2,0}$
such that
$\chi_{\mathrm{NP}}\cdot\left(E_{\psi}P_{V}-(\operatorname{Id}-G_{\psi})\right)\in{}^{3\mathrm{sc}}\Psi^{-\infty,-\infty}.$
The analogous statement holds near $\mathrm{SP}$, see Remark 5.7.
The lemma gives estimates:
Corollary 5.13.
Let $s,\ell\in\mathbb{R}$, $\psi\in C_{c}^{\infty}(\mathbb{R})$ with $\psi(s)=1$ for $|s|<\delta$. Let $Q,G\in{}^{3\mathrm{sc}}\Psi^{0,0}$. If ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(\chi_{\mathrm{NP}})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$, then
$\lVert Q(\operatorname{Id}-G_{\psi})u\rVert_{s+2,\ell}\lesssim\|GP_{V}u\|_{s,\ell}+\lVert u\rVert_{-N,-M},$
with the left hand side being finite if the right hand side is finite.
The analogous statement holds near $\mathrm{SP}$, see Remark 5.7.
Proof of corollary from lemma.
This follows exactly as in elliptic estimates, namely from Lemma 5.12, we have
$\lVert Q(\operatorname{Id}-G_{\psi})u\rVert_{s+2,\ell}$ $\lesssim\lVert Q\chi_{\mathrm{NP}}(\operatorname{Id}-G_{\psi})u\rVert_{s+2,\ell}+\lVert u\rVert_{-N,-M}$
$\lesssim\lVert Q\chi_{\mathrm{NP}}E_{\psi}P_{V}u\rVert_{s+2,\ell}+\lVert u\rVert_{-N,-M}$
$\lesssim\lVert GP_{V}u\rVert_{s,\ell}+\lVert u\rVert_{-N,-M}$
where the last line uses the boundedness properties of $Q\chi_{\mathrm{NP}}E_{\psi}\in{}^{3\mathrm{sc}}\Psi^{-2,0}$. ∎
Now we prove the lemma.
Proof.
It is more convenient to work with $F_{\psi}$ as its symbols will be defined using functional calculus of bounded operators. Thus we seek $F_{\psi}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ with
$\chi_{\mathrm{NP}}\cdot\left(F_{\psi}\circ(D_{t}^{2}+H_{V_{+}}+E)^{-1}P_{V}-(\operatorname{Id}-G_{\psi})\right)\in{}^{3\mathrm{sc}}\Psi^{-\infty,-\infty}.$
The lemma follows immediately by defining $E_{\psi}$ as in (5.13).
Now we construct $F_{\psi}$. We do so iteratively at the symbolic level. We will use repeatedly that if $\psi,\tilde{\psi}\in C_{c}^{\infty}(\mathbb{R})$,
$\psi\tilde{\psi}=\tilde{\psi}\implies(\operatorname{Id}-G_{\psi})(\operatorname{Id}-G_{\tilde{\psi}})=\operatorname{Id}-G_{\psi}.$
At each stage of the construction we shrink the region in which the symbol conditions are satisfied. Fixing $K\subset\{s:\psi(s)=1\}$, we take $\psi_{1}\in C_{c}^{\infty}(\mathbb{R})$ with $\operatorname{supp}\psi_{1}\subset K$. We take $F_{1}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ with
$E_{1}\coloneqq F_{1}(D_{t}^{2}+H_{V_{+}}+E)^{-1}P_{V}-(\operatorname{Id}-G_{\psi_{1}})\in{}^{3\mathrm{sc}}\Psi^{-1,-1}.$
The front face symbol condition is then exactly given by the function $f_{\psi_{1}}$ (5.13),
(5.14) $\hat{N}_{\operatorname{ff}}(F_{1})(\tau)=f_{\psi_{1}}((\tau^{2}+H_{V_{+}}+E)^{-1}(\tau^{2}-H_{V_{+}})).$
(That such an indicial family exists follows from an analogous construction in the $\mathrm{scl},\mathrm{sc}$ calculus.) The two other symbol conditions are easy to arrange as they can be solved on off the characteristic set of $P_{V}$. For the inductive step, assume that for some $\psi_{n}\in C_{c}^{\infty}(\mathbb{R})$ which is $1$ near 00, with $\operatorname{supp}\psi_{n}\subset K$, and $F_{1},\dots,F_{n}$ with $F_{j}\in{}^{3\mathrm{sc}}\Psi^{1-j,1-j}$ such that
$E_{n}\coloneqq(\operatorname{Id}-G_{\psi_{n}})(\sum_{j=1}^{n}F_{j})(D_{t}^{2}+H_{V_{+}}+E)^{-1}(D_{t}^{2}-H_{V_{+}})-(\operatorname{Id}-G_{\psi_{n}})\in\Psi^{-n,-n}.$
Then we solve for $F_{n+1}$ such that
$F_{n+1}(D_{t}^{2}+H_{V_{+}}+E)^{-1}(D_{t}^{2}-H_{V_{+}})+(\operatorname{Id}-G_{\psi_{n}})K_{n}\in{}^{3\mathrm{sc}}\Psi^{-n-1,-n-1}.$
which is to say $\hat{N}_{\operatorname{ff},-n}(F_{n+1})(\tau)=F_{\psi}((\tau^{2}+H_{V_{+}}+E)^{-1}(-\tau^{2}+H_{V_{+}}))\hat{N}_{\operatorname{ff},-n}(E_{n}).$ Then for any $\psi_{n+1}\in C_{c}^{\infty}(\mathbb{R})$ with $\psi_{n+1}\psi_{n}=\psi_{n}$,
$(\operatorname{Id}-G_{\psi_{n+1}})(\sum_{j=1}^{n+1}F_{j})(D_{t}^{2}+H_{V_{+}}+E)^{-1}(D_{t}^{2}-H_{V_{+}})-(\operatorname{Id}-G_{\psi_{n+1}})\in\Psi^{-n-1,-n-1}.$
Thus by induction there are $F_{1},\dots$, with $F_{j}\in{}^{3\mathrm{sc}}\Psi^{1-j,1-j}$ such that
$(\operatorname{Id}-G_{\psi})(\sum_{j=1}^{n}F_{j})(D_{t}^{2}+H_{V_{+}}+E)^{-1}(D_{t}^{2}-H_{V_{+}})-(\operatorname{Id}-G_{\psi})\in\Psi^{-n,-n},$
for any $n$, and taking the Borel sum gives
$F_{\psi}=(\operatorname{Id}-G_{\psi})(\sum_{j=1}^{\infty}F_{j})$
It remains only to know that $f_{\psi}((\tau^{2}+H_{V_{+}}+E)^{-1}(\tau^{2}-H_{V_{+}}))\in\Psi_{\mathrm{scl},\mathrm{sc}}^{0,0,0}$, but this follows from an identical parametrix argument. ∎
@@ 5.3. Localization near the characteristic set
While, (5.17) will ensure that (5.15) lies in ${}^{3\mathrm{sc}}\Psi^{0,0}$, the condition (5.16) will ensure that commutators with $P_{V}$ have the correct order (i.e. lose one order compared with composition.) This is established in the following proposition, in which we use $\rho=1/\tau$ to modify the differential order so that the centrality condition is preserved; this leads to the appearance of factors of $\langle\tau,\zeta\rangle/\tau$ in the local components of the symbol. Also (see Remark 2.4) we use powers of $x$ to rescale the symbol. Here we also define all symbols using a fiber defining function $\rho_{fib}$ which is equal to $\rho$ near the characteristic set.
This proof follows from the same arguments as in Vasy [36, Proposition 13.1]. The main difference is that since $G_{\psi}$ is of differential order 00 and vanishes to infinite order at $\operatorname{fibeq}$ (by Proposition 5.8), the product is a $3\mathrm{sc}$-operator if $q$ is smooth on $[{}^{\mathrm{sc}}T^{*}X;\operatorname{fibeq}]$.
@@
Let $q\in C^{\infty}([{}^{\mathrm{sc}}T^{*}X;\operatorname{fibeq}])$ satisfying (5.16), and assume that there is $c>0$ such that $\operatorname{supp}q\cap[-c,c]=\varnothing$. Then for $s,\ell\in\mathbb{R}$,
@@ 6. Propagation estimates over $C$
Let $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ and $V_{0}\in S^{-1}(\mathbb{R}^{n}_{z})$ such that $V-V_{0}\in{}^{3\mathrm{sc}}\Psi^{1,-1}$ and
@@
Let $B_{0}\in{}^{3\mathrm{sc}}\Psi^{s,\ell},E_{0}\in{}^{3\mathrm{sc}}\Psi^{s,\ell},G\in{}^{3\mathrm{sc}}\Psi^{s-1,\ell+1}$ for some $s,\ell\in\mathbb{R}$. Assume that
@@
    $\operatorname{WF}^{\prime}_{\operatorname{ff}}(E_{0})=\varnothing$,
@@
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0})\cup{}^{3\mathrm{sc}}\!\operatorname{Ell}(E_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$
@@
    For all $\alpha\in\gamma_{3\mathrm{sc}}({}^{3\mathrm{sc}}\!\operatorname{Ell}(G))$, we have that ${}^{\mathrm{sc}}\!H_{p}(\alpha)\not=0$.
@@
    For all $\alpha\in\operatorname{Char}(P_{0})$ that are incoming to ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0})$, in the sense that
    $\pi_{X,\tau}(\alpha)\in\pi_{X,\tau}(\gamma_{3\mathrm{sc}}({}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0}))\cap\operatorname{Char}(P_{0}))\,,$
    there exists $s_{\alpha}\in(-\varepsilon,0)$ such that
    $\exp(s_{\alpha}\,{}^{\mathrm{sc}}\!H_{p})(\alpha)\in{}^{3\mathrm{sc}}\!\operatorname{Ell}(E_{0})$
    and for all $s\in[s_{\alpha},0]$,
    $\exp(s\,{}^{\mathrm{sc}}\!H_{p})(\alpha)\in\gamma_{3\mathrm{sc}}({}^{3\mathrm{sc}}\!\operatorname{Ell}(G))\,.$
For each $u\in H_{\mathrm{sc}}^{-N,-M}$ with $E_{0}u\in L^{2}$, $GP_{V}u\in L^{2}$ it follows that $B_{0}u\in L^{2}$ and
$\lVert B_{0}u\rVert\lesssim_{M,N}\lVert E_{0}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
@@
In words, the proposition states that, to obtain estimates at a given $\tau$ level over $\mathrm{NP}$, we must control the backward flow out of the entire sphere $|\zeta|^{2}=\tau^{2}+m^{2}$ in a neighborhood of $\mathrm{NP}$. The elliptic set of $E_{0}$ must control this set in the sense that it must contain a transversal of the sphere earlier along the flow. In particular, the elliptic set of $G$ over $\operatorname{ff}$ must contain the whole of $\operatorname{WF}^{\prime}_{\operatorname{ff}}(B_{0})$.
Remark 6.3.
The same proposition holds if the wavefront set of $B_{0}$ is controlled by the for backward flow of the elliptic set of $E_{0}$, in which case $s_{\alpha}\in(0,\epsilon)$.
@@
$=2(1-M\delta\varrho)f^{\flat}f\cdot\hat{N}_{\operatorname{ff}}(G_{\psi})^{2}+G_{\psi}R_{\varphi}G_{\psi}$
@@
(6.24) $B$ $=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}b)G_{\phi}\,,$
(6.25) $E$ $=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}e)G_{\phi}\,,$
(6.26) $Q_{0}$ $=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}q)G_{\phi}\,.$
@@
There exists $F\in{}^{3\mathrm{sc}}\Psi^{2s-1,2\ell-1}$ and $\varepsilon\in(0,1)$ such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(F)\subset{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)$ and
$i[P_{V},G_{\phi}Q^{*}QG_{\phi}]+(M-2(2\ell+1))Q_{0}^{*}Q_{0}\leq-(1-\varepsilon)B^{*}B+E^{*}E+F\,.$
@@
$\hat{N}_{\operatorname{ff},\ell+1/2-\delta_{2}}(Q_{r})$ $=\rho^{\delta_{1}}\hat{N}_{\operatorname{ff},\ell+1/2}(Q)\,,$
$\hat{N}_{\operatorname{mf}}(Q_{r})$ $=\rho^{\delta_{1}}x^{\delta_{2}}\hat{N}_{\operatorname{mf}}(Q)\,,$
$\sigma_{3\mathrm{sc}}(Q_{r})$ $=\rho^{\delta_{1}}x^{\delta_{2}}\sigma_{3\mathrm{sc}}(Q)$
@@
There exists $F_{r}^{\prime}\in{}^{3\mathrm{sc}}\Psi^{2s-1-2\delta_{1},2\ell-1-2\delta_{2}}$ such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(F_{r}^{\prime})\subset{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)$ and
@@
We first prove a variant of the propagation estimate with the specific $B$ and $E$.
@@
Let $u\in H_{\mathrm{sc}}^{-N,-M}$, $B,E$ as in (6.24)-(6.25), and $G\in{}^{3\mathrm{sc}}\Psi^{s-1,\ell+1}$ with
${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\cup{}^{3\mathrm{sc}}\!\operatorname{Ell}(E)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G).$
If $Eu,GP_{V}u\in L^{2}$, then $Bu\in L^{2}$ and we have the estimate
$\lVert Bu\rVert\lesssim_{N,M}\lVert Eu\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
Remark 6.12.
Note that here we make no explicit assumption about the relationship between the Hamiltonian flow and the elliptic set of $G$ as in Proposition 6.1. This is because the operators here are given explicitly in (6.24)-(6.25). Here ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)$ by construction is an open neighborhood of $+\infty\in\overline{W^{\perp}}\subset\dot{C}_{3\mathrm{sc}}[X;C]$.
@@
We take $\delta_{1}=s+N$ and $\delta_{2}=\ell+M$. Then for $r\in(0,1)$, we have that

### v2

@@ 5.1. Functional calculus
This definition of $F_{\varphi}$ makes sense as a bounded operator on $L^{2}$ by continuous functional calculus since $(D_{t}^{2}+H_{V_{+}}+E)^{-1}P_{V_{+}}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ is a bounded symmetric operator. Note that $f_{\varphi}$ is not compactly supported, so Proposition 5.1 does not apply directly to $f(A)$ with $A\in{}^{3\mathrm{sc}}\Psi^{0,0}$. However, given an almost analytic extension $\tilde{\varphi}$ of $\varphi$, the function $\tilde{f}_{\varphi}(z)=(1-\tilde{\varphi}(z))/z$ is an almost analytic extension of $f_{\varphi}$. Since $|\tilde{f}_{\varphi}(z)|\lesssim 1/\langle z\rangle$, the Helffer-Sjöstrand formula (5.4) and thus Proposition 5.1 hold with $\varphi$ replaced by $f_{\varphi}$.
Proposition 5.14 is now a consequence of the mapping properties of $E_{\varphi}$, applied to $P_{V}u$. The “one order better” estimate from the operator $G^{\prime}$ is obtained by writing $P_{V}=P_{V_{+}}+V-V_{+}$ and using the assumptions on $V-V_{+}$. If $P_{V}$ itself admitted a parametrix with error $G_{\varphi}$, we would obtain an estimate with purely residual terms, i.e. no $G^{\prime}$ present, but this is not to be expected for $V$ which are not purely static for large $t$.
Proof of Proposition 5.14.
Let $E_{\varphi}\in{}^{3\mathrm{sc}}\Psi^{-2,0}$ be as in (5.13). We write
$B=BG_{\varphi}+B(\operatorname{Id}-G_{\varphi})=BG_{\varphi}+\tilde{B}P_{V}+R\,,$
where $\tilde{B}=BE_{\varphi}\in{}^{3\mathrm{sc}}\Psi^{-2,0}$ and $R=BE_{\varphi}(P_{V}-P_{V_{+}})\in{}^{3\mathrm{sc}}\Psi^{-1,-r}$. Then using Theorem 4.20, we obtain
$\lVert Bu\rVert_{s,\ell}\lesssim\lVert BG_{\varphi}u\rVert_{s,\ell}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-1,\ell-r}+\lVert u\rVert_{-N,-M}\,,$
which is what we wanted. ∎
@@ 5.2. Localization near the characteristic set
While, (5.16) will ensure that (5.14) lies in ${}^{3\mathrm{sc}}\Psi^{0,0}$, the condition (5.15) will ensure that commutators with $P_{V}$ have the correct order (i.e. lose one order compared with composition.) This is established in the following proposition, in which we use $\rho=1/\tau$ to modify the differential order so that the centrality condition is preserved; this leads to the appearance of factors of $\langle\tau,\zeta\rangle/\tau$ in the local components of the symbol. Also (see Remark 2.4) we use powers of $x$ to rescale the symbol. Here we also define all symbols using a fiber defining function $\rho_{\operatorname{fib}}$ which is equal to $\rho$ near the characteristic set.
This proof follows from the same arguments as in Vasy [36, Proposition 13.1]. The main difference is that since $G_{\psi}$ is of differential order 00 and vanishes to infinite order at $\operatorname{fibeq}$ (by Proposition 5.9), the product is a $3\mathrm{sc}$-operator if $q$ is smooth on $[{}^{\mathrm{sc}}\overline{T}^{*}X;\operatorname{fibeq}]$.
@@
Let $q\in C^{\infty}([{}^{\mathrm{sc}}\overline{T}^{*}X;\operatorname{fibeq}])$ satisfying (5.15). Then for $s,\ell\in\mathbb{R}$,
@@
Remark 5.16.
Note that we implicitly assume that
$x^{-\ell}\rho^{-s}q\in\langle t,z\rangle^{-\ell}\langle\tau,\zeta\rangle^{-s}C^{\infty}([{}^{\mathrm{sc}}\overline{T}^{*}X;\operatorname{fibeq}])\,.$
This condition is for instance satisfied if $q$ vanishes for $\lvert\tau\rvert$ small and supported near $\mathrm{NP}$.
While $\operatorname{Op}_{L}(q)$ is not in the $3\mathrm{sc}$-calculus in general, we can find $Q\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that $QG_{\psi}=\operatorname{Op}_{L}(q)G_{\psi}$.
Lemma 5.17.
Let $q\in C^{\infty}([{}^{\mathrm{sc}}\overline{T}^{*}X;\operatorname{fibeq}])$ satisfying (5.15) and $s,\ell\in\mathbb{R}$. Assume that there is $\tau_{0}\in\overline{W^{\perp}}$ such that $f(\tau_{0})\not=0$. Then there exists $Q\in{}^{3\mathrm{sc}}\Psi^{s,\ell}$ such that
$QG_{\psi}=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}q)G_{\psi}$
and $\tau_{0}\in{}^{3\mathrm{sc}}\!\operatorname{Ell}(Q)$.
Proof.
Choose $\varphi\in C_{c}^{\infty}(\mathbb{R})$ with $\varphi|_{\operatorname{supp}\psi}\equiv 0$ so that $\varphi\psi=\psi$. We set
$Q=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}q)G_{\varphi}+\tilde{Q}(\operatorname{Id}-G_{\varphi})\,,$
where $\tilde{Q}\in{}^{\mathrm{sc}}\Psi^{s,\ell}$ with $\hat{N}_{\operatorname{ff}}(\tilde{Q})=f(\tau_{0})\operatorname{Id}$. This $Q$ satisfies the claimed equality, since $G_{\varphi}G_{\psi}=G_{\psi}$ and $\hat{N}_{\operatorname{ff},\ell}(Q)(\tau_{0})=\tau_{0}^{s}f(\tau_{0})\operatorname{Id}$, and hence $\tau_{0}\in{}^{3\mathrm{sc}}\!\operatorname{Ell}(Q)$. ∎
@@ 6. Propagation estimates over $C$
We introduce the notion of control under the Hamilton flow that arises in the propagation estimates.
Definition 6.1.
Let $U,V,W\subset\dot{C}_{3\mathrm{sc}}[X;C]$. We say that $U$ is (backward) controlled by $V$ through $W$ if for all $\alpha\in\operatorname{Char}(P_{0})$ that are incoming to $U$, in sense that
$\pi_{X,\tau}(\alpha)\in\pi_{X,\tau}(\gamma_{3\mathrm{sc}}(U)\cap\operatorname{Char}(P_{0}))\,,$
there exists $s_{\alpha}<0$ such that
$\exp(s_{\alpha}{}^{\mathrm{sc}}\!H_{p})(\alpha)\in V$
and for all $s\in[s_{\alpha},0]$,
$\exp(s{}^{\mathrm{sc}}\!H_{p})(\alpha)\in W\,.$
The following is the main propagation estimate we use near $C$ and away from the radial sets.
@@
Let $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ be asymptotically static of order $r\geq 1$ and
@@
Moreover, we assume that $H_{V_{\pm}}$ have purely absolutely continuous spectrum in $[m^{2},\infty)$.
Let $\delta>0$ sufficiently small, $\varphi,\psi_{1},\psi_{2}\in C_{c}^{\infty}(\mathbb{R})$ with $\operatorname{supp}\varphi\subset(-\delta,\delta)$ and $\psi_{j}|_{(-\delta,\delta)}\equiv 1$, and $B,E,G,G^{\prime},B^{\prime}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that
@@
    $\operatorname{WF}_{\operatorname{ff}}^{\prime}(E)=\varnothing$,
@@
    ${}^{\mathrm{sc}}\!H_{p}(\alpha)\not=0$ for all $\alpha\in\gamma_{3\mathrm{sc}}({}^{3\mathrm{sc}}\!\operatorname{Ell}(G))$,
@@
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})$ is controlled by ${}^{3\mathrm{sc}}\!\operatorname{Ell}(E)$ through ${}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$,
@@
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})$.
For all $M,N,s,\ell\in\mathbb{R}$ and $u\in H_{\mathrm{sc}}^{-N,-M}$ with $EG_{\psi_{1}}u\in H_{\mathrm{sc}}^{s,\ell}$, $GG_{\psi_{2}}P_{V}u\in H_{\mathrm{sc}}^{s-1,\ell+1}$, $G^{\prime}P_{V}u\in H_{\mathrm{sc}}^{s-2,\ell}$, and $B^{\prime}u\in H_{\mathrm{sc}}^{s-1,\ell-r}$, it follows that $Bu\in H_{\mathrm{sc}}^{s,\ell}$ and
$\lVert Bu\rVert_{s,\ell}$ $\leq C\big{(}\lVert EG_{\psi_{1}}u\rVert_{s,\ell}+\lVert GG_{\psi_{2}}P_{V}u\rVert_{s-1,\ell+1}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-1,\ell-r}$
[…]
@@
In words, the proposition states that, to obtain estimates at a given $\tau$ level over $\mathrm{NP}$, we must control the backward flow out of the entire sphere $|\zeta|^{2}=\tau^{2}+m^{2}$ in a neighborhood of $\mathrm{NP}$. The elliptic set of $E_{0}$ must control this set in the sense that it must contain a transversal of the sphere earlier along the flow. In particular, the elliptic set of $G$ over $\operatorname{ff}$ must contain the whole of $\operatorname{WF}^{\prime}_{\operatorname{ff}}(BG_{\varphi})$.
@@
Remark 6.4.
Since the estimates are microlocal, it suffices to prove the case $V_{\pm}=V_{0}$ for some $V_{0}\in{}^{3\mathrm{sc}}\Psi^{1,0}$ and
$V-V_{0}\in{}^{3\mathrm{sc}}\Psi^{1,-r}\,.$
@@
$=2(1-M\delta\varrho)f^{\flat}f\cdot\hat{N}_{\operatorname{ff}}(G_{\psi})^{2}+\hat{N}_{\operatorname{ff}}(G_{\psi})R_{\varphi}\hat{N}_{\operatorname{ff}}(G_{\psi})$
@@
(6.24) $B_{0}$ $=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}b)G_{\phi}\,,$
(6.25) $E_{0}$ $=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}e)G_{\phi}\,,$
(6.26) $Q_{0}$ $=\operatorname{Op}_{L}(x^{-(\ell+1/2)}\rho^{-(s-1/2)}q)G_{\phi}\,,$
(6.27) $Q_{0}^{\prime}$ $=\operatorname{Op}_{L}(x^{-\ell}\rho^{-s}q)G_{\phi}\,.$
@@
Let $O\subset\dot{C}_{3\mathrm{sc}}[X;C]$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\subset O$. There exists $F\in{}^{3\mathrm{sc}}\Psi^{2s-1,2\ell-1}$ and $\varepsilon\in(0,1)$ such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(F)\subset O$ and
$i[P_{V},Q_{0}^{*}Q_{0}]+(M-2(2\ell+1))(Q_{0}^{\prime})^{*}Q_{0}^{\prime}\leq-(1-\varepsilon)B^{*}B+E^{*}E+F\,.$
@@
$\hat{N}_{\operatorname{ff},\ell+1/2-\delta_{2}}(Q_{r})$ $=\rho^{\delta_{1}}\hat{N}_{\operatorname{ff},\ell+1/2}(Q_{0})\,,$
$\hat{N}_{\operatorname{mf}}(Q_{r})$ $=\rho^{\delta_{1}}x^{\delta_{2}}\hat{N}_{\operatorname{mf}}(Q_{0})\,,$
$\sigma_{3\mathrm{sc}}(Q_{r})$ $=\rho^{\delta_{1}}x^{\delta_{2}}\sigma_{3\mathrm{sc}}(Q_{0})$
@@
Let $O\subset\dot{C}_{3\mathrm{sc}}[X;C]$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\subset O$. There exists $F_{r}^{\prime}\in{}^{3\mathrm{sc}}\Psi^{2s-1-2\delta_{1},2\ell-1-2\delta_{2}}$ such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(F_{r}^{\prime})\subset O$ and
@@
We first prove a variant of the propagation estimate with the specific $B_{0}$, $Q_{0}$, and $E_{0}$.
@@
With $B_{0}$, $Q_{0}$ and $E_{0}$ defined as above, and $B_{1}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})$. If $u\in H_{\mathrm{sc}}^{-N,-M}$ with $B_{1}u\in H_{\mathrm{sc}}^{s-1/2,\ell-1/2}$, $Q_{0}P_{V}u\in H_{\mathrm{sc}}^{-1/2,1/2}$, and $E_{0}u\in L^{2}$, then $B_{0}u\in L^{2}$ and
$\lVert B_{0}u\rVert\leq C\left(\lVert Q_{0}P_{V}u\rVert_{-1/2,1/2}+\lVert E_{0}u\rVert+\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}+\lVert u\rVert_{-N,-M}\right)\,.$
@@
Using the operators and parameters defined above, we take $\delta_{1}=s+N$ and $\delta_{2}=\ell+M$. Then for $r\in(0,1)$, we have that